hi how to do 6 and 7) pleaseYou are interested in taking out a loan of $30,000 from your local bank. The bank charges you an interest rate of 5% per year
Answer #1

HI

6) This question is an example of annuity problem

where Present Value P = $30,000

interest rate r = 5%

time t = 10 years

periodic payment C =?

Present value of annuity P = (C/r)*(1-(1+r)^-n)

30000 = (C/0.05)*(1-1.05^-10)

30000*0.05 = C*(1-0.614)

C = 1500/0.386

C = $3,885.14

hence option C is correct here.

7) Profitability index = incremental profit from segment / number of constraint resource required

hence the higher the profitability, better will be the project.

Hence option D is false here.

Thanks